# A Survey on Multilingual Large Language Models: Corpora, Alignment, and Bias

### v2

@@ V-B Bias Benchmark
This section focuses on the issue of bias evaluation in MLLMs. Extensive studies have developed varied datasets and approaches that serve as benchmarks for bias assessment. In this section, we provide a thorough review of these benchmarks. Table VI illustrates benchmarks commonly used for evaluating bias. Notably, these datasets primarily focus on bias attributes related to gender and occupation [146, 147, 148], predominantly available in English [149, 150, 151, 152]. Several datasets also encompass languages such as Spanish, German, and French [128] [134].
@@
Based on the tasks and languages, benchmarks in Table VI can be categorized into three types: general benchmarks, task-specific benchmarks, and language-specific benchmarks.
@@
General benchmarks mainly refer to evaluation benchmarks that have a wide range of applications and can be used for different tasks, including some major evaluation metrics and datasets. For example, Association Test (WEAT, SEAT, and CEAT) [153, 154, 155] are widely used to measure bias in word-, sentence- and contextualized-level embeddings; GLUE [150] is designed to measure the impact that the introduced debiasing techniques will have on downstream performance by evaluating the capabilities of the NLP model.
@@
Multilingual benchmarks refer to the benchmark datasets in multilingual contexts, including MIBs [128] and MozArt [134]. The lack of robust multilingual evaluation benchmarks poses significant barriers to assessing biases in multilingual contexts. Therefore, creating more multilingual evaluation datasets is an urgent problem to be solved. One potential solution is to translate existing bias benchmarks that mainly only cover English [156, 157]. Nevertheless, it is important to note that translated benchmarks may introduce additional biases due to translation errors and cultural differences. Thus, when designing a multilingual bias benchmark, it’s crucial to consider various cultural contexts and develop cultural-diverse datasets [12].
| Benchmark | Time | Benchmark/Metric | Type of bias | Goal |
| General Category | ||||
| WEAT [153] | 2017 | Evaluation metric | Gender | Measure bias in word embeddings |
| GLUE [150] | 2018 | English benchmark | Untargeted | Evaluate how debiasing techniques affect downstream task performance. |
| SEAT [154] | 2019 | Evaluation metric | Gender | Measure bias in sentence encoders. |
| CEAT [155] | 2020 | Evaluation metric | Untargeted | Measure bias in contextualized word embeddings. |
| InBias [128] | 2020 | Evaluation metric | Gender, Occupation | Quantify intrinsic bias in multilingual word embeddings. |
| ExBias [158] | 2020 | Evaluation metric | Gender, Occupation | Measure debiasing word embeddings by comparing their performance before and after debiasing. |
| StereoSet [14] | 2020 | English benchmark | Gender, Occupation, Race etc. | Evaluate the stereotypical biases of popular PLMs. |
| Task-specific Category | ||||
| Winogender [146] | 2018 | English benchmark | Gender, Occupation | Identify bias in in coreference resolution systems. |
| WinoBias [147] | 2018 | English benchmark | Gender, Occupation | Identify bias in coreference resolution systems. |
| EEC [149] | 2018 | English benchmark | Gender, Race | Measure bias of race and gender through differences in predicting sentiment intensity between sentences. |
| CrowS-Pairs [151] | 2018 | English benchmark | Race, Religion, Age etc. | Measure certain social bias in LLMs. |
| WinoMT [148] | 2019 | English benchmark | Gender | Investigate gender bias in machine translation systems. |
| BiosBias [152] | 2019 | English benchmark | Gender, Occupation | Evaluate bias in predicting individual occupation based on their short biography. |
| FairFace [159] | 2019 | Face Attribute benchmark | Gender, Race, Age | Evaluate how to mitigate bias in existing databases by collecting more diverse facial images. |
| Language-specific Category | ||||
| MIBs [128] | 2020 | English, Spanish, German, and French benchmark | Gender, Occupation | Conduct the intrinsic bias analysis. |
| MozArt [134] | 2022 | English, Spanish, German, and French benchmark | Gender, Language | Evaluate whether MLLMs are equally fair to demographic groups across languages. |
V-C Debias Technique
@@
V-C1 Model Debiasing
The existing methods for debiasing models can be categorized into four lines: representation based methods, pre-training based methods, fine-tuning based methods, and prompt based methods.
@@ V-C1 Model Debiasing
Word-level methods:They focus on static [161] or contextual embedding representations [162]. For example, INLP [161] was proposed to remove bias like race, gender, and age in static word embeddings with iterative null-space projection-based debiasing method. Linguistic Identity Removal (LIR) [162] was proposed to addressed bias in multilingual contextual word embeddings. It utilized singular value decomposition and orthogonal projection to identify and remove linguistic information in multilingual semantic space.
Pre-training based methods. In this approach, debiasing occurs during the pre-training stage, where the parameters of LLMs are modified to align with fairness criteria such as SEAT [154]. Dropout as proposed in [163], is a bias mitigation technique using dropout regularization [164]. By adjusting dropout parameters in BERT and ALBERT for attention weights and hidden activations, along with performing an extra phrase of pre-training, gender bias within these models can be alleviated. However, this method cannot guarantee whether the bias associations may resurge when the debiased models are fine-tuned on downstream tasks [165].
@@
For example, Guo et al. [169] proposed a framework named Auto-Debias, using cloze-style prompts to probe, identify and correct the biases in PLMs. This method first searches for the biased prompts probes the biased content with such prompts, and then corrects the model bias. Mattern et al. [170] explored GPT-3’s stereotypical associations with genders and jobs and proposed a framework to quantify and further reduce these biases using debiasing prompts. They also discussed prompt selection with varying degrees of abstraction and concluded that more concrete debiasing prompts exhibited a more pronounced effect. Dhingra et al. [171] demonstrated that employing a method involving chain-of-thought prompting through SHAP analysis can efficiently mitigate biases against queer people in the output of LLMs. Schick et al. [172] introduced a debiasing technique named Self-Debias which uses a model’s internal knowledge to discourage biased text generation. It starts by utilizing hand-crafted prompts to encourage the model to generate toxic text. Subsequently, a second continuation that is non-discriminatory can be produced from the model by scaling down the probabilities of tokens considered likely under the first toxic generation.
V-C2 Data Debiasing
@@
Language bias mitigation. Language bias in MLLMs is caused by the imbalanced language proportion, acting as the dominance of English and other major languages in the available multilingual training corpora. Constructing more balanced corpora has proven to be an effective solution for mitigating language bias. For example, XNLI [173] was developed to support 15 languages on the evaluation of XLU, providing information-rich standard evaluation tasks for cross-language sentence understanding. In addition, the release of CulturaX [174], a multilingual dataset that includes 167 languages and a total of 63,000 tokens, addresses the lack of open-source and easy-to-use datasets for effectively training multilingual large models. Furthermore, The ROOTS dataset [175] was developed to cover 59 languages with a total size of 1.6TB.
@@
Demographic bias mitigation. Demographic bias occurs when data overly emphasizes or represents a certain specific population. The commonly used method for mitigating demographic bias is counterfactual data augmentation. Based on identifying biased terms, it creates text that contradicts existing facts, reducing over-reliance on specific scenarios or groups and mitigating biases stemming from class imbalances within data. With the method, model’s reliance on false features can largely reduce and thus enhance the model’s robustness. Counterfactual augmented data is mainly achieved through two methods: manual generation and model generation, where both of them achieved comparable quality of generation [177]. Existing studies [178, 179, 180] have shown that counterfactual data augmentation is a simple and effective approach to mitigate bias in data.
@@
As a conclusion, although some progress has been made in data debiasing, there are still some limitations and challenges. First, although language bias can be mitigated by constructing more balanced corpora, how to effectively collect and integrate high-quality low-resource language texts is still a problem [176]. Second, although counterfactual data augmentation alleviates demographic bias to some extent, it may affect other performance of the model. Therefore, how to make debiased data prevent model generating bias while maintaining model performance is a promising direction worth exploring in depth.
VI Future Directions
@@
Ethical Impact of Multilingual LLMs. Multilingual LLMs can inherit biases present in their training data, leading to ethical risk of generation. Due to the high proportion of Western language data in training data, the MLLMs are inclined to reflect Western-centric concepts [190]. How to mitigate biases and ensure fairness and culturally sensitive in text generation are key challenges for the further development of MLLMs.

### v3

@@ 5.2 Bias Benchmark
This section focuses on the issue of bias evaluation in MLLMs. Extensive studies have developed varied datasets and approaches that serve as benchmarks for bias assessment. In this section, we provide a thorough review of these benchmarks. Table 7 illustrates benchmarks commonly used for evaluating bias. Notably, these datasets primarily focus on bias attributes related to gender and occupation [153, 154, 157], predominantly available in English [155, 149, 156, 158]. Several datasets also encompass languages such as Spanish, German, and French [130] [136].
@@
Based on the tasks and languages, benchmarks in Table 7 can be categorized into three types: general benchmarks, task-specific benchmarks, and language-specific benchmarks.
@@
General benchmarks mainly refer to evaluation benchmarks that have a wide range of applications and can be used for different tasks, including some major evaluation metrics and datasets. For example, Association Tests (WEAT, SEAT, and CEAT) [148, 150, 151] are widely used to measure bias in word-, sentence-, and contextualized-level embeddings; GLUE [149] is designed to measure the impact that the introduced debiasing techniques will have on downstream performance by evaluating the capabilities of the NLP model.
@@
Multilingual benchmarks refer to the benchmark datasets in multilingual contexts, including MIBs [130] and MozArt [136]. The lack of robust multilingual evaluation benchmarks poses significant barriers to assessing biases in multilingual contexts. Therefore, creating more multilingual evaluation datasets is an urgent problem to be solved. One potential solution is to translate existing bias benchmarks that mainly only cover English [160, 161]. Nevertheless, it is important to note that translated benchmarks may introduce additional biases due to translation errors and cultural differences. Thus, when designing a multilingual bias benchmark, it’s crucial to consider various cultural contexts and develop culturally diverse datasets [12].
5.3 Debias Technique
Bias in MLLMs rises significant ethical concerns, potentially leading to serious consequences. Demographic bias, in particular, can result in the unfairly representation or treatment of certain groups, thereby perpetuating societal inequalities. For instance, if gender bias is present in reference letters generated by MLLMs and not properly addressed, it could harm the success rates of female applicants [162]. Similarly, language bias can reinforce cultural stereotypes and misunderstandings, inadvertently exacerbating negative perceptions against other cultures when models favor certain languages or cultural contexts, thereby undermining efforts to promote inclusivity and diversity. Additionally, evaluation bias compromises the reliability and fairness of model evaluations, leading to skewed performance metrics and misinformed decisions. Mitigating these biases is crucial for ensuring ethical integrity, fairness, and transparency in MLLM applications.
@@ 5.3 Debias Technique
5.3.1 Model Debiasing
The existing methods for debiasing models can be categorized into four lines according to their debiasing stage: representation based methods, pre-training based methods, fine-tuning based methods, and prompt based methods.
@@ 5.3.1 Model Debiasing
Word-level methods:They focus on static [164] or contextual embedding representations [165]. For example, INLP [164] was proposed to remove bias like race, gender, and age in static word embeddings with iterative null-space projection-based debiasing method. Linguistic Identity Removal (LIR) [165] was proposed to address bias in multilingual contextual word embeddings. It utilizes singular value decomposition and orthogonal projection to identify and remove linguistic information in multilingual semantic space.
Pre-training based methods. In this approach, debiasing occurs during the pre-training stage, where the parameters of LLMs are modified to align with fairness criteria such as SEAT [150]. Dropout as proposed in [166], is a bias mitigation technique using dropout regularization [167]. By adjusting dropout parameters in BERT and ALBERT for attention weights and hidden activations, along with performing an extra phase of pre-training, gender bias within these models can be alleviated. However, this method cannot guarantee whether the bias associations may resurge when the debiased models are fine-tuned on downstream tasks [168].
@@
For example, Guo et al. [172] proposed a framework named Auto-Debias, using cloze-style prompts to probe, identify, and correct the biases in PLMs. This method first searches for the biased prompts, probes the biased content with such prompts, and then corrects the model bias. Mattern et al. [173] explored GPT-3’s stereotypical associations with genders and jobs and proposed a framework to quantify and further reduce these biases using debiasing prompts. They also discussed prompt selection with varying degrees of abstraction and concluded that more concrete debiasing prompts exhibited a more pronounced effect. Dhingra et al. [174] demonstrated that employing a method involving chain-of-thought prompting through SHAP analysis can efficiently mitigate biases against queer people in the output of LLMs. Schick et al. [175] introduced a debiasing technique named Self-Debias which uses a model’s internal knowledge to discourage biased text generation. It starts by utilizing hand-crafted prompts to encourage the model to generate toxic text. Subsequently, a second continuation that is non-discriminatory can be produced from the model by scaling down the probabilities of tokens considered likely under the first toxic generation.
5.3.2 Data Debiasing
@@
Language bias mitigation. Language bias in MLLMs is caused by the imbalanced language proportion, acting as the dominance of English and other major languages in the available multilingual training corpora. Constructing more balanced corpora has proven to be an effective solution for mitigating language bias. For example, XNLI [176] was developed to support 15 languages on the evaluation of XLU, providing information-rich standard evaluation tasks for cross-language sentence understanding. In addition, the release of CulturaX [177], a multilingual dataset that includes 167 languages and a total of 63,000 tokens, addresses the lack of open-source and easy-to-use datasets for effectively training multilingual large models. Furthermore, the ROOTS dataset [178] was developed to cover 59 languages, with a total size of 1.6TB.
@@
Demographic bias mitigation. Demographic bias occurs when data overly emphasizes or represents a certain specific population. The commonly used method for mitigating demographic bias is counterfactual data augmentation. Based on identifying biased terms, it creates text that contradicts existing facts, reducing over-reliance on specific scenarios or groups and mitigating biases stemming from class imbalances within data. With the method, model’s reliance on false features can be largely reduced, thereby enhancing the model’s robustness. Counterfactual augmented data is mainly achieved through two methods: manual generation and model generation, both of which achieve comparable quality of generation [180]. Existing studies [181, 182, 183] have shown that counterfactual data augmentation is a simple and effective approach to mitigate bias in data.
@@
5.3.3 Effectiveness and Challenges
Debiasing techniques for MLLMs have shown promise in mitigating biases and improving fairness. Both model and data debiasing approaches play vital roles in addressing ethical concerns, with complementary effectiveness and challenges. Effectiveness from different techniques can be summarized as follow:
Flexibility and Adaptability from Model Debiasing. By directly modifying the model’s internal representations or outputs, model debiasing allows for targeted adjustments to address specific tasks and biases. These techniques have been shown to significantly reduce stereotypes in text generation [175, 173] and question answering [164, 169], achieving quantifiable improvements in fairness metrics while maintaining competitive task performance. This approach is advantageous as it does not require changes to the training data, making it both efficient and adaptable. However, its impact is often limited when foundational biases within the training data remain unaddressed.
Systemic and Long-term Improvements from Data Debiasing. Data debiasing alleviate biases at their source by improving the quality and balance of the training data used to pre-train MLLMs. This approach excels in addressing systemic issues such as language imbalances and demographic biases, enabling MLLMs to perform more fairly across both high- and low-resource languages, as well as different demographic groups. This is demonstrated by improved accuracy in cross-lingual understanding [2, 5] and enhanced fairness in demographic sentiment analysis tasks [181, 183].
Despite progress in addressing ethical concerns through debiasing, challenges still persist in achieving comprehensive and scalable solutions.
Trade-offs with Performance. Many debiasing techniques introduce trade-offs, where mitigating bias can result in reduced accuracy or fluency in downstream tasks [154, 187]. Balancing fairness and performance remains a critical concern, especially in low-resource languages, where available data is often scarce and the risk of model overfitting or bias amplification is higher. This challenge is a promising direction worth exploring in depth.
Dependency on Context. The inherent diversity of language and culture makes achieving absolute fairness across all contexts difficult [188]. While constructing more balanced corpora is a viable solution for mitigating language bias, the effective collection and integration of high-quality, low-resource language texts remains a formidable challenge [179], often requiring significant human, financial, and computational resources.
Lack of Interpretability. Existing methods for bias understanding and mitigation face a major challenge: the lack of interpretability. Biases may stem not only from the model itself but also from external factors like training data, algorithms, or task settings. Research into bias interpretability helps identify these sources, enabling more targeted debiasing strategies. Methods like bias attribution in neural networks [169] and identifying bias neurons [189] are crucial for improving fairness and transparency.
6 Future Directions
@@
Ethical Impact of Multilingual LLMs. Multilingual LLMs can inherit biases present in their training data, leading to ethical risks of generation. Due to the high proportion of Western language data in training data, the MLLMs are inclined to reflect Western-centric concepts [196]. How to mitigate biases and ensure fairness and cultural sensitivity in text generation are key challenges for the further development of MLLMs.